# Quantum Complexity in Constrained Many-Body Models: Scars, Fragmentation, and Chaos

### v1

@@ Acknowledgements.
S.S.R. wants to thank Germán Sierra for introducing the QGL model and sharing valuable comments on the manuscript. We also want to thank Philipp Hauke, Himadri Shekhar Dhar, Anindita Bera for reading the manuscript and sharing useful comments. We thank Soumik Bandopadhyay for useful discussion on spectral form factor. S.S.R. acknowledges the faculty research scheme at IIT (ISM) Dhanbad, India under Project No. FRS/2024/PHYSICS/MISC0110. We also acknowledge the use of the QuSpin package [63] for the numerical simulations presented in this work.
@@ Appendix A Krylov subspace structures of $H^{\mathcal{N}(0)}$
Appendix B Description of unfolding process of energy spectrum
Figure 11: (a) The staircase function $\widetilde{N}(E)$ and corresponding fitted polynomial N(E){N}(E) for $H_{\text{QGL}}$. (b) Shows similar plot for $H_{\text{Tot}}$. We can see the imprint of the localization kind of property of $H_{\text{Tot}}$ is also visible here. Because of this particular plot, we had to use a high-degree polynomial $(15\text{th order})$ for better fitting, which we used in all other Hamiltonians to maintain uniformity.
@@ Appendix B Description of unfolding process of energy spectrum
Appendix C Spin-flip symmetry resolutions for $H^{\mathcal{N}(2)}$

### v2

@@ Acknowledgements.
S.S.R. wants to thank Germán Sierra for introducing the QGL model and sharing valuable comments on the manuscript. We also want to thank Philipp Hauke, Himadri Shekhar Dhar, Anindita Bera for reading the manuscript and sharing useful comments. We thank Soumik Bandyopadhyay for useful discussion on spectral form factor. S.S.R. acknowledges the faculty research scheme at IIT (ISM) Dhanbad, India under Project No. FRS/2024/PHYSICS/MISC0110. We also acknowledge the use of the QuSpin package [70] for the numerical simulations presented in this work.
@@ Appendix A Krylov subspace structures of $H^{\mathcal{N}(0)}$
Appendix B Fragmentation scaling of different models
Figure 16: Scaling of biggest Krylov subspace in fragmented models. Shown is the ratio $d_{L}/\mathcal{D}_{L}$ as a function of system size $L$ for (a) $H^{\mathcal{N}(0)}$, (b) $H^{\mathcal{N}(2)}$, (c) $H^{\mathrm{Pert}}_{\mathrm{PPXPP}}(k=1)$, and (d) $H^{\mathrm{Pert}}_{\mathrm{PPXPP}}(k=2)$.
We can see that the Hamiltonian $H^{\mathcal{N}(0)}$ exhibits strong fragmentation. But it contains many Krylov subspaces of dimension $1$, i.e. frozen states, leading to a strongly suppressed $d_{L}/\mathcal{D}_{L}$. $H^{\mathcal{N}(2)}$ shows a clear even-odd effect in its scaling with $L$. All three models shown in (a), (b) and (c) in Fig. 16 show that $d_{L}/\mathcal{D}_{L}$ decreases as L increases. However, the model $H_{\text{PPXPP}}^{\text{Pert}}(k=2)$ (shown in (d)) shows that the ratio $d_{l}/\mathcal{D}_{L}$ is saturating around 0.5 with increase in L, giving a qualitatively different behavior than others. This demonstrates that Krylov subspace growth is strongly influenced by both kinetic constraints and perturbations, leading to qualitatively different fragmentation patterns.
Appendix C Description of unfolding process of energy spectrum
Figure 17: (a) The staircase function $\widetilde{N}(E)$ and corresponding fitted polynomial N(E){N}(E) for $H_{\text{QGL}}$. (b) Shows similar plot for $H^{\text{Tot}}$. We can see the imprint of the localization kind of property of $H^{\text{Tot}}$ is also visible here. Because of this particular plot, we had to use a high-degree polynomial $(15\text{th order})$ for better fitting, which we used in all other Hamiltonians to maintain uniformity.
@@ Appendix C Description of unfolding process of energy spectrum
Appendix D Spin-flip symmetry resolutions for $H^{\mathcal{N}(2)}$